# Batch speculative decoding Done right

### v1

@@ 3.1 Minimal Batch Realignment: EqSpec
Algorithm 3 EXSpec: Cross-Batch Scheduling
1: Draft and Target model $\mathcal{M}_{d}$, $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Window size $W$, Batch size $B$
2: Generated sequences $\mathcal{S}$
3: $\textit{Pool}\leftarrow\texttt{InitSequencePool}(\mathcal{P})$
4: $\triangleright$ Tokenize and optionally sort by length
5: $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
6: while Pool.hasActive() do
7:   Phase 1: Lazy Realignment
8:   $\mathcal{B},\textit{mask},\textit{KV}\leftarrow\texttt{GetBatch}(\textit{Window},B)$
9:   $\triangleright$ Try same-length concatenation
10:   $\triangleright$ Fallback to Unpad-Repad Realignment
11:   Phase 2: Draft Generation
12:   $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{B},\textit{mask},K)$
13:   Phase 3: Batch Verification
14:   $A,B,\textit{KV}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{B},D,\textit{KV})$
15:   Phase 4: Write-Back and Window Refill
16:   for $i\in\mathcal{B}$ do
17:    $\textit{Pool}[i]\leftarrow\textit{Pool}[i]\oplus A[i]\oplus B[i]$
18:    $\textit{Pool.KV}[i]\leftarrow\textit{KV}[i]$
19:    if $\textit{isComplete}(\textit{Pool}[i])$ then $\textit{Pool.deactivate}(i)$
21: return Pool.sequences
Two observations follow. First, when sequences within a batch share the same length, batch realignment overhead is negligible; likewise, with very small draft models or prompt lookahead decoding (Fu et al., 2024b), $c_{\text{draft}}$ becomes small relative to verification. Second, and critically, $c_{\text{overhead}}(B)$ scales superlinearly with batch size $B$ due to the ragged-tensor problem (Section 4.3). This overhead decomposes as $c_{\text{overhead}}(B)=c_{\text{pad}}(B)+c_{\text{kv}}(B)$, where $c_{\text{pad}}(B)$ accounts for unpad–append–repad operations and $c_{\text{kv}}(B)$ for KV-cache realignment on rank-4 tensors. While $c_{\text{draft}}$ and $c_{\text{verify}}$ benefit from GPU parallelism and remain relatively stable within the hardware’s batch-parallel regime, the alignment overhead grows with both batch size and variance in acceptance rates across sequences.
3.2 Cross-Batch Scheduling: EXSpec
Profiling EqSpec shows that alignment overhead consumes 39.4% of computation at batch size 8, rising to 46.7% at batch size 16. Because this cost is inherent to synchronizing ragged tensors, micro-optimizing the primitives yields limited gains. Instead of accelerating these operations, EXSpec avoids them by scheduling.
@@ 3.2 Cross-Batch Scheduling: EXSpec
EXSpec differs from EqSpec in how it manages sequence lifecycles. Rather than maintaining fixed batches that require realignment after every verification, we introduce a SequencePool that holds sequences individually in their ragged states. This enables three optimizations: (i) sequences that complete (reach EOS) are immediately removed, avoiding wasted computation on finished items; (ii) lazy realignment defers synchronization until strictly necessary, keeping sequences in their natural ragged form between steps; and (iii) dynamic batch formation over a sliding window of $W>B$ sequences greatly increases the chance of finding same-length groups that can be concatenated without any realignment.
@@
Figure 4 illustrates the flow. After verification, ragged sequences return directly to the pool without realignment. The scheduler then scans the active window and attempts to form batches of identical length. Same-length sequences concatenate directly—no padding adjustments, no position-ID recomputation, no KV-cache realignment—bypassing $c_{\text{overhead}}(B)$. Only when same-length grouping fails do we fall back to the expensive unpad–append–repad procedure from EqSpec. Algorithm 3 formalizes this in four phases: dynamic batch formation (prioritizing same-length groups), draft generation, verification, and pool write-back with continuous window refresh. Combined with prompt-length sorting, grouping rates approach unity for similar workloads, turning the worst case (constant realignment) into the rare case.
@@ 4 Experiments
We evaluate EqSpec and EXSpec along two dimensions often conflated in prior work: correctness (whether outputs match non-speculative generation) and throughput (tokens per second). Through systematic evaluation across three model families and comparisons with both research prototypes and production systems, we show that our approach uniquely preserves output equivalence while achieving competitive speedups.
@@ 4.1 Experimental Setup
We use SpecBench (Xia et al., 2024), focusing on the first turn of each conversation because our evaluation targets offline batch inference rather than multi-turn interaction. We also use Multi30k (Elliott et al., 2016) for a controlled EXSpec study that contrasts random sampling with an identical-length subset, isolating sequence-length diversity as the driver of grouping rate. For the main evaluation, we measure: (1) Throughput: tokens/s across batch sizes; and (2) Correctness: exact-match rate (full-sequence equivalence with non-speculative decoding) and partial-match rate (fraction of tokens matching until the first divergence). The partial-match metric helps localize failure modes—early divergence typically indicates position-ID or KV-cache misalignment.
@@
Following our design-space taxonomy (Section 2), we evaluate: (1) Masking approaches: BSP (Su et al., 2023) attempts masking with adaptive speculation but suffers position-ID inconsistencies (BASS (Qian et al., 2024) also follows this approach but requires custom CUDA kernels, limiting generality); (2) Dynamic-padding approaches: DSD (Yan et al., 2025) explores padding but mishandles the KV-cache, while EqSpec implements correct synchronization and EXSpec adds cross-batch scheduling; (3) Reference baselines: Spec-1 (batch-size-1 speculation from Hugging Face Transformers), which does not support batch speculative decoding [^2] https://github.com/huggingface/transformers/issues/32165. We also compare with production systems vLLM [^3] We use the vLLM v0 engine because v1 deprecates speculative decoding. (Kwon et al., 2023) (which subsumes TETRIS (Wu et al., 2025) and TurboSpec (Liu et al., 2025) as vLLM forks) and SGLang-EAGLE [^4] SGLang is compatible only with the EAGLE family; we compare Vicuna-7B/EAGLE2 and Qwen3-8B/EAGLE3. There are no available weights for GLM-4. (Zheng et al., 2024; Li et al., 2024b; 2025), noting that their continuous batching and memory-management designs complicate direct comparison. We further test SGLang-EAGLE-Deterministic (SGLang Team, 2025), which enables deterministic execution to reduce numerical variance. No existing implementation uses rollback due to its inherent wastefulness.
@@ E P E P E P E P E P E P
Continuous Batching Systems (vs. own non-spec)
vLLM + Spec 96.9 / 98.0 65.6 / 78.5 72.7 / 84.5
SGLang + EAGLE 69.8 / 79.5 47.7 / 65.4
SGLang + EAGLE + Det 85.0 / 90.0 50.6 / 69.5
Table 1: Correctness check reports exact match (E) and partial match (P) scores. Top: compares with batch=1 non-spec baseline. Bottom: with each system’s non-speculative mode (scores reported as E / P). Our approach sustains $95\%$ accuracy, while prior work (DSD, BSP) suffers major drops from KV-cache and position ID errors.
Figure 5: Decomposing batch speculative decoding performance. (a) Batch scaling efficiency normalized to BS=1, isolating GPU parallelism from per-sequence speculation; EXSpec initially exceeds the No-Ragged-Scaling upper bound before degrading due to alignment overhead. (b) Alignment overhead grows super-linearly with batch size, consuming up to 38% of inference time, validating that $c_{\text{overhead}}(B)$ dominates at scale. (c) Cross-batch grouping rates on Multi30k for random vs. uniform-length sequences, showing that length homogeneity transforms grouping effectiveness.
4.2 Output Correctness Verification
We verify correctness using deterministic greedy decoding to eliminate sampling variance and enable precise bug isolation. This avoids metrics such as ROUGE (Qian et al., 2024), which can mask implementation failures (e.g., repetitive corruption can still score reasonably). Instead, we use exact match (any divergence) and partial match (fraction of tokens before the first mismatch) to diagnose failure modes.
Table 1 reveals distinct patterns. Our methods maintain $\approx$95% exact match across settings, the remaining $\approx$5% divergence stems from numerical non-determinism in floating-point operations (He & Lab, 2025) and tie-breaking in argmax sampling rather than algorithmic errors. By contrast DSD and BSP fail catastrophically with different signatures. DSD’s near-zero scores indicate immediate position-ID misalignment—the model fails from the first token. BSP shows higher partial match (up to 39.7%) but low exact match, indicating gradual degradation: outputs are initially correct before KV-cache drift misdirects attention, triggering repetition. These complementary patterns—immediate failure vs. gradual decay—show how partial-match metrics reveal not just that implementations fail, but when and why.
Production systems (vLLM, SGLang) introduce additional complexity. While both are accurate on Vicuna, they degrade markedly on Qwen3. SGLang-EAGLE-Deterministic helps disambiguate the cause: improved accuracy suggests most divergences stem from floating-point non-determinism (He & Lab, 2025) rather than algorithmic bugs. Despite this, we show below that their throughput still falls below non-speculative baselines.
@@ 4.3 Overhead and Scaling Dynamics
Figure 1 shows that EXSpec outperforms EqSpec in absolute throughput by combining speculative and batching gains, yet EqSpec exhibits negative scaling beyond BS=8. To test whether batch speculative decoding can retain GPU-parallelism benefits despite raggedness, Figure 5(a) normalizes each method’s throughput to its BS=1 baseline, isolating batch-level scaling from per-sequence speculation and enabling comparison to No-Ragged-Scaling (standard batched decoding without speculation). A notable effect emerges: EXSpec initially surpasses even this upper bound because speculation converts memory-bound operations into compute-bound verification, improving GPU utilization when managed correctly. However, at larger batch sizes, alignment overhead dominates and the advantage inverts, confirming that $c_{\text{overhead}}(B)$ eventually overwhelms parallelism gains.
@@
Production systems limitation.
Figure 6 shows that production frameworks struggle with batch speculative decoding. Although our method is not directly comparable to continuous-batching systems (which vary effective batch size dynamically), the trend is consistent: vLLM’s speculative decoding underperforms its baseline at high concurrency (leading to deprecation in v1), and SGLang+EAGLE is slower than non-speculative generation across all batch sizes. Two architectural factors likely contribute: (i) speculation repurposes the prefill stage for parallel verification, breaking the prefill–decode separation these systems optimize; and (ii) continuous batching already induces raggedness from varying request lengths, and speculation compounds this with per-sequence acceptance variance, creating nested raggedness that overwhelms existing schedulers. These outcomes indicate that speculation cannot simply be grafted onto architectures optimized for predictable token-by-token decoding; deeper redesign is required.
Method TPS $|\text{Spec}|$ Time/Verif. (ms) Verif. % Draft % Overhead %
EqSpec 95.6 1469 24.88 44.9 27.4 27.7
EXSpec 156.4 952 29.13 55.9 29.5 14.6
Table 2: Overhead anatomy of batch speculation methods. Despite slower per-verification time, EXSpec achieves higher throughput by reducing total verification calls and minimizing alignment overhead through intelligent scheduling.
@@ 5 Related Work
Speculative decoding accelerates LLM inference by verifying draft tokens in parallel. Two verification paradigms dominate: sequence verification—as in SpecDec and follow-ups (Xia et al., 2023; Santilli et al., 2023; Yang et al., 2023; Hooper et al., 2023; Zhang et al., 2023; Fu et al., 2024a)—which preserves the target model’s output distribution; and tree verification—e.g., Medusa/EAGLE variants (Miao et al., 2024; Spector & Re, 2023; Sun et al., 2023; He et al., 2023; Cai et al., 2024; Li et al., 2024a; b; 2025)—which explores multiple branches to raise acceptance rates. Recent works parallelize multiple draft sequences per request (Stewart et al., 2024; Lee et al., 2024) but still verify each request independently. Approximate schemes (Kim et al., 2023; Zhong et al., 2025) trade exactness for speed; our work assumes lossless verification to detect implementation errors rather than approximation artifacts.
@@
Extending speculative decoding to batch settings introduces the ragged tensor problem: when sequences accept different numbers of draft tokens, the resulting variable-length tensors break GPU-friendly rectangular operations (Qian et al., 2024). Existing approaches face fundamental limitations detailed in Section 2. Position ID/masking approaches like BSP (Su et al., 2023) suffer from position ID inconsistencies across iterations. Dynamic padding approaches reveal critical implementation errors: both DSD (Yan et al., 2025) and Meta’s recent work (Tang et al., 2025) incorrectly sample bonus tokens from the draft model’s distribution rather than the target model’s, violating the fundamental correctness guarantee of speculative decoding. This error compounds with improper KV-cache handling, producing corrupted outputs—BSP generates repetitive tokens while DSD produces <unk> symbols. BASS (Qian et al., 2024) sidesteps these issues through custom CUDA kernels but sacrifices portability. These failures demonstrate that the ragged tensor problem extends beyond performance to correctness itself, motivating our cross-batch scheduling approach that maintains correctness while achieving batch scaling without custom kernels.
Production Systems for Speculative Decoding.
Serving stacks integrate speculation atop general batching (vLLM, SGLang/EAGLE, and forks such as TETRIS and TurboSpec (Kwon et al., 2023; Zheng et al., 2024; Li et al., 2025; Wu et al., 2025; Liu et al., 2025)). In practice, variable draft acceptance clashes with continuous batching and paged attention, and our measurements indicate scaling limits (e.g., small single-sequence gains that reverse at higher concurrency); some engines have since reduced or deprecated speculative-decoding support. We intentionally build our reference implementation from scratch, prioritizing correctness checks; compatibility with continuous batching, paged attention, and prefill–decode separation is orthogonal and left as future work.
@@ 6 Conclusion
We presented a correctness-first approach to batch speculative decoding that resolves the fundamental tension between maintaining output equivalence and achieving practical speedups. By identifying the precise synchronization invariants required for correctness, we explained why existing methods fail and established the minimal requirements for valid computation. In particular, our EqSpec implementation enforces these invariants but reveals that realignment overhead consumes up to 40% of computation. By contrast, EXSpec overcomes this limitation via cross-batch scheduling that dynamically groups same-length sequences to bypass realignment entirely. Moreover, experiments across three model families show that our approach uniquely preserves >>95% output equivalence while achieving up to 3× throughput improvement at batch size 8, hereby validating that a careful co-design of correctness constraints and scheduling can multiply batch parallelism with per-sequence speculation gains. Our findings on ragged shape tensor synchronization overhead and cross-batch scheduling may inform future designs for production systems, where integrating speculation with continuous batching remains an open challenge.
Acknowledgment
This work is supported by an eBay research grant.
Appendix A Appendix

### v2

@@ 3.2 EXSpec: Reducing Overhead via Cross-Batch Scheduling
Critically, $c_{\text{overhead}}(B)$ scales superlinearly with batch size $B$. Inside the alignment overhead, the KV-cache is the major bottleneck, consisting of rank-4 tensors (batch $\times$ heads $\times$ sequence $\times$ dimension), and each padding adjustment triggers allocation and concatenation of high-dimensional zeros. While $c_{\text{draft}}$ and $c_{\text{verify}}$ benefit from GPU parallelism, the alignment overhead grows with both batch size and variance in acceptance rates across sequences. Profiling EXSpec confirms this: alignment consumes 39.4% of computation at batch size 8, rising to 46.7% at batch size 16.
However, the speedup analysis reveals a key insight: when all sequences accept the same number of tokens, $\delta_{i}=0$ and realignment overhead vanishes (Corollary B.3). Rather than accelerating these operations, EXSpec sidesteps them through scheduling.
Cross-Batch Scheduling.
@@
EXSpec differs from EqSpec in how it manages sequence lifecycles. Rather than maintaining fixed batches that require realignment after every verification, we introduce a SequencePool that holds sequences individually in their ragged states. This enables three optimizations: (i) sequences that complete (reach EOS) are immediately removed, avoiding wasted computation; (ii) lazy realignment defers synchronization until strictly necessary; and (iii) dynamic batch formation over a sliding window of $W>B$ sequences greatly increases the chance of finding same-length groups.
@@
Figure 4 illustrates the flow. After verification, ragged sequences return directly to the pool without realignment. The scheduler scans the active window and attempts to form batches of identical length. Same-length sequences concatenate directly—no padding adjustments, no position-ID recomputation, no KV-cache realignment—bypassing $c_{\text{overhead}}(B)$ entirely. Only when same-length grouping fails do we fall back to unpad–append–repad. Combined with prompt-length sorting, grouping rates approach unity for similar workloads, turning constant realignment into the rare case. The complete procedure is detailed in Algorithm 2.
@@ 4 Experiments
We evaluate EqSpec and EXSpec along two dimensions often conflated in prior work: output equivalence and throughput. Through systematic evaluation across three model families and comparisons with both research prototypes and production systems, we show that our approach uniquely preserves output equivalence while achieving competitive speedups.
@@ 4.1 Experimental Setup
We use SpecBench (Xia et al., 2024), a widely adopted benchmark for speculative decoding research. We also use Multi30k (Elliott et al., 2016) for a controlled EXSpec study that contrasts random sampling with an identical-length subset, isolating sequence-length diversity as the driver of grouping rate. For the main evaluation, we measure: (1) Throughput: tokens/s across batch sizes; and (2) Output Equivalence: exact-match rate (full-sequence equivalence with non-speculative decoding) and partial-match rate (fraction of tokens matching until the first divergence). The partial-match metric helps localize failure modes—early divergence typically indicates position-ID or KV-cache misalignment.
@@
Following our design-space taxonomy (Section 2), we evaluate: (1) Masking approaches: BSP (Su et al., 2023) attempts masking with adaptive speculation but suffers position-ID inconsistencies (BASS (Qian et al., 2024) also follows this approach but requires custom CUDA kernels, limiting generality); (2) Dynamic-padding approaches: DSD (Yan et al., 2025) explores padding but mishandles the KV-cache, while EqSpec implements correct synchronization and EXSpec adds cross-batch scheduling; (3) Reference baselines: Spec-1 (batch-size-1 speculation from Hugging Face Transformers), which does not support batch speculative decoding [^2] https://github.com/huggingface/transformers/issues/32165. Production systems (vLLM, SGLang; see also Sections 2 and 3.1) employ continuous batching with paged attention, which sidesteps the ragged tensor problem through different architectural choices. However, neither currently supports batch speculative decoding with external draft models: vLLM deprecated this in v1 due to batch expansion memory overhead, and SGLang only supports EAGLE-family drafters. We include a comparison in Appendix E for completeness, but note these systems solve a related but distinct problem. No existing implementation uses rollback due to its inherent wastefulness.
@@ E P E P E P E P E P E P
Table 1: Output Equivalence of speculative decoding implementations. Exact match (E) and partial match (P) scores compared against batch=1 non-speculative baseline. Our approach sustains $95\%$ accuracy, while prior work (DSD, BSP) suffers major drops from KV-cache and position ID errors.
Figure 5: Decomposing batch speculative decoding performance. (a) Batch scaling efficiency: each method’s throughput normalized to its own BS=1 baseline, isolating scaling behavior from absolute performance. The No-Ragged-Scaling line shows measured autoregressive decoding throughput (not a theoretical bound), serving as a reference for how standard batching scales without ragged tensor overhead. (b) Alignment overhead grows super-linearly with batch size, consuming up to 38% of inference time, validating that $c_{\text{overhead}}(B)$ dominates at scale. (c) Cross-batch grouping rates on Multi30k for random vs. uniform-length sequences, showing that length homogeneity transforms grouping effectiveness.
4.2 Output Equivalence Verification
We verify output equivalence using deterministic greedy decoding to eliminate sampling variance and enable precise bug isolation. This avoids metrics such as ROUGE (Qian et al., 2024), which can mask implementation failures (e.g., repetitive corruption can still score reasonably). Instead, we use exact match (any divergence) and partial match (fraction of tokens before the first mismatch) to diagnose failure modes.
Table 1 reveals distinct patterns. Our methods maintain $\approx$95% exact match across settings, the remaining $\approx$5% divergence stems from numerical non-determinism in floating-point operations (He and Lab, 2025; Gond et al., 2026) and tie-breaking in argmax sampling rather than algorithmic errors. This attribution is corroborated by Liu et al. (2025b), who observe that generation length varies significantly across batch sizes due to kernel-level nondeterminism even in functionally correct implementations. By contrast DSD and BSP fail catastrophically with different signatures. DSD’s near-zero scores indicate immediate position-ID misalignment—the model fails from the first token. BSP shows higher partial match (up to 39.7%) but low exact match, indicating gradual degradation: outputs are initially correct before KV-cache drift misdirects attention, triggering repetition. These complementary patterns—immediate failure vs. gradual decay—reflect distinct root causes: DSD suffers from batch-independent errors (wrong sampling distribution) compounded by synchronization failures, while BSP accumulates KV-cache drift from position-ID desynchronization. A complete taxonomy distinguishing batch-independent bugs from batch-specific synchronization errors appears in Appendix A. We provide additional analysis of production systems (vLLM, SGLang) in Appendix E.
@@ 4.3 Overhead and Scaling Dynamics
Figure 1 shows that EXSpec outperforms EqSpec in absolute throughput by combining speculative and batching gains, yet EqSpec exhibits negative scaling beyond BS=8. To test whether batch speculative decoding can retain GPU-parallelism benefits despite raggedness, Figure 5(a) measures batch scaling efficiency: each method’s throughput at batch size $N$ divided by its own throughput at batch size 1. This normalization isolates how well each method scales with batch size, independent of absolute performance differences. The No-Ragged-Scaling line represents actual measured autoregressive decoding throughput (without speculation), similarly normalized to its own BS=1 baseline; it is not a theoretical upper bound but rather a reference showing how standard batching scales when no ragged tensor synchronization is required. A notable effect emerges: EXSpec initially exceeds this reference line. This occurs because each method is normalized independently, and speculation converts memory-bound token generation into compute-bound verification, which benefits more from GPU parallelism. Although EXSpec may have lower absolute throughput than standard decoding at BS=1 due to draft model overhead, its relative scaling from BS=1 to BS=8 can surpass that of non-speculative decoding. However, at larger batch sizes, alignment overhead dominates and the advantage inverts, confirming that $c_{\text{overhead}}(B)$ eventually overwhelms parallelism gains. The key finding is that valid batch speculative decoding need not sacrifice batch scaling efficiency, given that prior methods either produce gibberish outputs or fail to scale.
@@
Method TPS $|\text{Spec}|$ ms/V V% D% OH%
EqSpec 95.6 1469 24.9 44.9 27.4 27.7
EXSpec 156.4 952 29.1 55.9 29.5 14.6
Table 2: Overhead anatomy of batch speculation methods. V%=verification, D%=draft, OH%=overhead. Despite slower per-verification time, EXSpec achieves higher throughput by reducing verification calls and alignment overhead.
@@ 5 Related Work
Speculative decoding accelerates LLM inference by verifying draft tokens in parallel. Two verification paradigms dominate: sequence verification—as in SpecDec and follow-ups (Xia et al., 2023; Santilli et al., 2023; Yang et al., 2023; Hooper et al., 2023; Zhang et al., 2023; Fu et al., 2024)—which preserves the target model’s output distribution; and tree verification—e.g., Medusa/EAGLE variants (Miao et al., 2024; Spector and Re, 2023; Sun et al., 2023; He et al., 2023; Cai et al., 2024; Li et al., 2024b, a, 2025)—which explores multiple branches to raise acceptance rates. Recent works parallelize multiple draft sequences per request (Stewart et al., 2024; Lee et al., 2024) but still verify each request independently. Approximate schemes (Kim et al., 2023; Zhong et al., 2025) trade exactness for speed; our work assumes lossless verification to detect implementation errors rather than approximation artifacts.
@@
Extending speculative decoding to batch settings introduces the ragged tensor problem: when sequences accept different numbers of draft tokens, the resulting variable-length tensors break GPU-friendly rectangular operations (Qian et al., 2024). Existing approaches face fundamental limitations detailed in Section 2. Position ID/masking approaches like BSP (Su et al., 2023) suffer from position ID inconsistencies across iterations. Dynamic padding approaches reveal critical implementation errors: both DSD (Yan et al., 2025) and Meta’s recent work (Tang et al., 2025) incorrectly sample bonus tokens from the draft model’s distribution rather than the target model’s, violating the fundamental output equivalence guarantee of speculative decoding. This error compounds with improper KV-cache handling, producing corrupted outputs—BSP generates repetitive tokens while DSD produces <unk> symbols. BASS (Qian et al., 2024) sidesteps these issues through custom CUDA kernels but sacrifices portability. Production systems (vLLM, SGLang) avoid the problem entirely via continuous batching but do not support external draft models at batch scale (Appendix E). These failures reveal that batch speculative decoding is fundamentally challenge—yet no prior work formalizes what invariants must hold. We address this gap by establishing synchronization requirements and providing the first correct implementation.
@@ 6 Conclusion
We establish that valid batch speculative decoding requires maintaining synchronization invariants, specifically rectangular alignment and position-ID contiguity, which prior implementations fail to preserve. Our theoretical analysis formalizes these invariants and proves that the EQSPEC update rule preserves them across verification rounds, while our cost analysis reveals that alignment overhead grows superlinearly with batch size as an inherent cost of algorithmic correctness. EXSPEC reduces this overhead by exploiting uniform acceptance rates, where realignment cost vanishes, through cross-batch scheduling that dynamically groups same-length sequences. Empirically, our methods achieve 95% exact match across three model families while delivering up to 3× throughput at batch size 8, demonstrating that the synergy between batching and speculation is achievable only through principled synchronization.
Impact Statement
This work focuses on improving the efficiency of LLM inference without altering model outputs or introducing biases. By providing the first valid batch speculative decoding implementation with output equivalence guarantees, we establish a reliable foundation for future research and production deployment. Prior implementations produced corrupted outputs, limiting their practical utility and potentially misleading subsequent work built upon them. Our approach ensures that acceleration techniques do not compromise model reliability. We will release our implementation as open source upon acceptance, providing the research community with a verified reference implementation for batch speculative decoding.
Appendix Contents
  • A
    Bug Taxonomy in Prior Implementations
  • B
    Proof of Synchronization Invariants
  • C
    BatchVerify Algorithm Details
  • D
    Latency-Throughput Tradeoff in Online Serving
  • E
    Continuous Batching Systems
  • F
    Discussion
  • G